OpenStudy (anonymous):

The circumference is 24pie and the measure of arc CD is 60 degrees. What is the length of CD?

OpenStudy (anonymous):



OpenStudy (anonymous):

Can you solve this?\[ \large \frac{24 \pi}{360^\circ} = \frac{x}{60^\circ} \]

OpenStudy (anonymous):

one moment...

OpenStudy (anonymous):

The angle will be proportional to the arc length. Just thing of the circumference as being the arc length of a \(360^\circ \) arc.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I think it should be \[\frac{ x }{ C } = \frac{ \angle }{ 360 }\]

OpenStudy (anonymous):

so \[\frac{ x }{ 24\pi } = \frac{ 60 }{ 360 }\] solve for x using that

OpenStudy (anonymous):

i don't have a graphing calculator, but: x/24pie=.17

OpenStudy (anonymous):

no, wait: would i cross multiply?

OpenStudy (anonymous):

x = (60/360) * 24pi = (1/6) * 24pi = 4pi

OpenStudy (anonymous):

After multiplying both sides by 24pi

OpenStudy (anonymous):

i see, so 4pi is the length of that arc (CD)

OpenStudy (anonymous):

RIGHT?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

THANKS!